HELP! I need assistance with this demand supply problem!
The demand and supply equations for a certain item are given by
D=-5p+40
S=-p^2+30p-8
Find the equilibrium price!
Thanks in advance!
ac

Answer by Edwin McCravy(20086)

HELP! I need assistance with this demand supply problem!
The demand and supply equations for a certain item are given by
D=-5p+40
S=-p^2+30p-8
Find the equilibrium price!
Thanks in advance!
ac

If the SUPPLY is more than the DEMAND, then there are some left 
over that go to waste.
If the DEMAND is more than the SUPPLY, then there is not enough 
to go around.

But when SUPPLY and DEMAND are equal, there is equilibrium.

So to find what price causes equilibrum we set the DEMAND equal 
to the SUPPLY.

            D = S

     -5p + 40 = -pē + 30p - 8

Can you solve that?


You'll have to use the quadratic formula.  There will be two
solutions, p = 1.42984132 and p = 33.57015868

You have to discard the larger because that value of p causes
the supply and demand to be a negative number.  So the 
equilibrium price rounded off to the nearest penny is
$1.43 if you are in the US, or pounds of euros if you are in
Europe.

Edwin